OpenStudy (anonymous):

i need help with algebra 2 please? ill medal:)

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (studygurl14):

\[\sqrt[4]{1296x^{16}y^{12}}\]

OpenStudy (anonymous):

@StudyGurl14 do you know what it would be simplified?

OpenStudy (studygurl14):

\[\sqrt[4]{(81 * 16)(x^{4}x^{4}x^{4}x^{4})(y^{4}y^{4}y^{4})}\]

OpenStudy (anonymous):

thats not an option..?

OpenStudy (studygurl14):

I'm not done

OpenStudy (anonymous):

ooh..lol ok:)

OpenStudy (anonymous):

@StudyGurl14 could you possibly hurry? i have to leave soon..:/

OpenStudy (studygurl14):

\[(4)(\left| x \right| \left| x \right| \left| x \right| \ \left| x \right|)(\left| y \right|\left| y \right|\left| y \right|)\sqrt[4]{81}\] \[4\left| x^{4} \right|\left| y^{3} \right|\sqrt[4]{9 * 9}\]

OpenStudy (studygurl14):

If you're in a hurry, it is A. btw

OpenStudy (anonymous):

oh, ok thank you very much:)

OpenStudy (anonymous):

mmm im not sure about that...

OpenStudy (studygurl14):

hold on! @Hailey553 I made a mistake. It isn't A!

OpenStudy (anonymous):

what is 6^4 and 36^4?

OpenStudy (anonymous):

haha! ok:)

OpenStudy (studygurl14):

\[\sqrt[4]{3^{4}} = 3\] so 3 *4 = 12 hm...I made amistake cause that's not an option

OpenStudy (anonymous):

try taking it backwards... instead of thinking of 4th root. Take everything to the 4th power.

OpenStudy (anonymous):

so, which one would it be?

OpenStudy (studygurl14):

Oh it is B or C. something with the 6

OpenStudy (anonymous):

its B if you have to leave right now

OpenStudy (anonymous):

but stay for explaination

OpenStudy (studygurl14):

I made mistake up there with the 16

OpenStudy (anonymous):

\[\sqrt[4]{1296x ^{16}y ^{12}}=\left( 6^{4}\left( x ^{4} \right)^{4} \left( y ^{3} \right)^{4}\right)^{\frac{ 1 }{ 4 }}\] \[=\left\{ \left( 6x ^{4}y ^{3} \right)^{4} \right\}^{\frac{ 1 }{4 }}=6x ^{4}y ^{3}\]

OpenStudy (anonymous):

nice job @surjithayer

OpenStudy (anonymous):

don't forget the absolute value signs though.

OpenStudy (anonymous):

thank you guys so much!

OpenStudy (studygurl14):

@surjithayer did it a much better way than I did @Hailey553. You should fan + medal him/her